error in measurement of radius of a cylinder  is 2% and in length it is 1%. the error in the measurement of the volume of cylinder is
 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
volume of cylinder = pi r^2 l
 
now differentiating wrt themselves
 
2 dr/r + dl/l = percentqge error in volume
 
 =  2*2 + 1 = 4 + 1 = 5%